a coffee company wants a new flavor of Cajun coffee. How many pounds of coffee worth $7 a pound should be added to 20 pounds of coffee worth $4 a pound to get a mixture worth $5 a pound?

Answer by richwmiller(17219)

7a+20*4=5*(a+20)
7a+80=5a+100
2a=20
a=10